# A Path Integral Treatment of Time-dependent Dunkl Quantum Mechanics

### v1

@@ Abstract
This paper presents an analytical treatment of the path integral formalism for time-dependent quantum systems within the framework of Wigner-Dunkl mechanics, emphasizing systems with varying masses and time-dependent potentials. By employing generalized canonical transformations, we reformulated the path integral to develop an explicit expression for the propagator. This formalism is applied to specific cases, including a Dunkl-harmonic oscillator with time-dependent mass and frequency. Solutions for the Dunkl-Caldirola-Kanai oscillator and a model with a strongly pulsating mass are derived, providing exact propagator expressions and corresponding wave functions. These findings extend the utility of Dunkl operators in quantum mechanics, offering new insights into the dynamics of time-dependent quantum systems.
@@ 1 Introduction
where denotes the Wigner parameter. In 1989, Dunkl introduced a new type of differential operator , defined as a linear combination of partial derivatives weighted by the reflection operator, allowing the study of special functions and their properties in a more general setting [29, 30]. The one-dimensional version of the Dunkl operator is given by
@@
Our aim in this paper is to explore the Dunkl path integral formalism for the propagator in the presence of a time-dependent potential. The treatment is mainly based on suitably chosen generalized canonical transformations and auxiliary equations. This paper is organized as follows: In Sec. 2, we reformulate the path integral formalism for the propagator of the time-dependent system. In Sec. 3, we consider the case of the Dunkl-harmonic oscillator with a time-dependent mass and frequency, and we calculate the relative propagator and the wave functions. In Sec. 4, we illustrate the general procedure by considering two models of varying mass. Finally, Sec. 5 contains the conclusion.
@@ 2 Time-dependent propagator
We start by outlining a method to compute the path integral for non-relativistic Wigner-Dunkl quantum systems with a time-dependent mass in the presence of general time-dependent potentials. This approach employs explicitly time-dependent spacetime transformations to facilitate the calculations, with the systeme’s dynamics governed by a Wigner-Dunkl Hamiltonian:
@@
To enhance the manageability of the path integral and facilitate its calculations, it is useful to utilize the method of canonical transformations [17], specifically by performing a time-dependent canonical transformation defined as,
@@
We now decompose the kernel propagator into two parity components, even and odd, distinguished by their respective eigenvalues, . The propagator for the even parity case is given by:
@@ 3 Harmonic oscillator
where is the modified Bessel function. Returning to the original variable and applying the decomposition outlined in Eq. (5), we can express the even and odd kernel propagators as follows:
@@ 5 Conclusion
Time-dependent quantum mechanics has become a pivotal area of research due to its applications across diverse fields, including plasma physics, quantum optics, and quantum chemistry. While the path integral method is widely used in time-dependent quantum systems, its application within the Dunkl formalism remains largely unexplored. In this study, we extend the path integral approach to time-dependent Dunkl quantum mechanics by reducing it to a stationary form through explicit generalized canonical and time transformations.
More precisely, we derived a general expression for the propagator and developed explicit formulations for the Dunkl harmonic oscillator, with a particular focus on two analytically solvable cases involving harmonic oscillators with time-dependent mass. This framework was applied to the Dunkl-harmonic oscillator with time-dependent mass and frequency, enabling the calculation of exact propagators and corresponding wave functions. These results underscore the applicability of Dunkl operators within time-dependent quantum systems, highlighting their versatility in addressing non-stationary dynamics. Subsequently, the effectiveness of this approach was further validated by its application to two illustrative models: the Dunkl-Caldirola-Kanai oscillator, characterized by an exponentially varying mass, and a harmonic oscillator with a strongly pulsating mass. These examples provided concrete instances of how Dunkl mechanics can effectively accommodate complex dependencies of mass and frequency, thereby expanding the toolkit for analyzing time-dependent quantum systems.
Overall, this study demonstrates the capacity of the Dunkl formalism to address complex time-dependent quantum problems and derive exact solutions in cases that pose significant analytical challenges. This approach broadens the scope of Dunkl mechanics for applications in both theoretical and applied physics, offering novel insights into systems where time-dependent characteristics and reflection symmetries play a critical role.
@@ Acknowledgments
B. C. L. is grateful to Excellence Project PřF UHK 2211/2023-2024 for the financial support.

### v2

@@ Abstract
This study addresses a significant gap in the literature by extending the path integral formalism to time-dependent systems within the Wigner-Dunkl framework, deriving exact propagator solutions for analytically solvable cases. By employing generalized canonical transformations, we reformulated the path integral to develop an explicit expression for the propagator. This formalism is applied to specific cases, including a Dunkl-harmonic oscillator with time-dependent mass and frequency. Solutions for the Dunkl-Caldirola-Kanai oscillator and a model with a strongly pulsating mass are derived, providing exact propagator expressions and corresponding wave functions. These findings extend the utility of Dunkl operators in quantum mechanics, offering new insights into the dynamics of time-dependent quantum systems and possibly find application in quantum optics, plasma physics, and other fields.
@@ 1 Introduction
where denotes the Wigner parameter. In 1989, Dunkl introduced a new type of differential operator, the Dunkl operator, , defined as a linear combination of partial derivatives weighted by the reflection operator, allowing the study of special functions and their properties in a more general setting [29, 30]. The one-dimensional version of the Dunkl operator is given by
@@
Our aim in this paper is to explore the Dunkl path integral formalism for the propagator in the presence of a time-dependent potential. Unlike other approaches, such as the Lewis-Riesenfeld invariant method, the Dunkl formalism naturally integrates reflection symmetries, making it particularly suitable for systems with time-dependent potential and mass. The treatment is mainly based on suitably chosen generalized canonical transformations and auxiliary equations. This paper is organized as follows: In Sec. 2, we reformulate the path integral formalism for the propagator of the time-dependent system. In Sec. 3, we consider the case of the Dunkl-harmonic oscillator with a time-dependent mass and frequency, and we calculate the relative propagator and the wave functions. In Sec. 4, we illustrate the general procedure by considering two models of varying mass. Finally, Sec. 5 contains the conclusion.
@@ 2 Time-dependent propagator
We start by outlining a method to compute the path integral for non-relativistic Wigner-Dunkl quantum systems with a time-dependent mass in the presence of general time-dependent potentials. The dynamics of the system are governed by the Wigner-Dunkl Hamiltonian, which incorporates time-dependent mass and potential terms as follows:
@@
Our purpose here is to transform the time-dependent mass and potential terms into a stationary form using a set of transformations. We employ the method of canonical transformations [17], specifically performing a time-dependent canonical transformation defined as,
@@
We now decompose the kernel propagator into two parity components: the even kernel () and the odd kernel (), distinguished by their respective eigenvalues, . The propagator for the even parity case is given by:
@@ 3 Harmonic oscillator
where is the modified Bessel function. Here, we have to underscore that the Dunkl approach extends classical propagators for harmonic oscillators by incorporating reflection symmetries, leading to solutions involving modified Bessel functions, instead of the standard trigonometric or exponential functions for the standard case, that provide parity-specific insights into the system’s dynamics.
Returning to the original variable and applying the decomposition outlined in Eq. (5), we can express the even and odd kernel propagators as follows:
@@ 4.2 The harmonic oscillator with strongly pulsating mass
The results for the Dunkl-Caldirola-Kanai oscillator reveal that the exponentially varying mass effectively models quantum damping effects, highlighting the system’s relevance to dissipative quantum mechanics, while the harmonic oscillator with a strongly pulsating mass captures periodic energy fluctuations, providing insights into systems with time-dependent interactions such as cavity quantum electrodynamics.
4.3 Comparison of Dunkl Formalism and Traditional Approaches
Before concluding the manuscript, it is important to highlight the key distinctions between the Dunkl formalism and traditional methods for propagators. These differences are summarized in the following table:
| Aspect | Traditional Approach | Dunkl Formalism |
| Symmetry Consideration | Parity symmetry not explicitly integrated | Incorporates reflection symmetry explicitly |
| Mathematical Tools | Standard derivatives and trigonometric/exponential functions | Dunkl operator and modified Bessel, Legendre and Laguerre functions |
| Propagator Decomposition | Not applicable | Even and odd parity components |
| Applications | General quantum systems | Systems with parity/reflection symmetries |
@@ 5 Conclusion
Time-dependent quantum mechanics has become a pivotal area of research due to its applications across diverse fields, including plasma physics, quantum optics, and quantum chemistry. While the path integral method is widely used in time-dependent quantum systems, its application within the Dunkl formalism remains largely unexplored. In this study, we extended the path integral approach to time-dependent Dunkl quantum mechanics by reducing it to a stationary form through explicit generalized canonical and time transformations.
More precisely, we derived a general expression for the propagator and developed explicit formulations for the Dunkl harmonic oscillator. A particular focus was placed on two analytically solvable cases involving harmonic oscillators with time-dependent mass: the Dunkl-Caldirola-Kanai oscillator and a harmonic oscillator with a strongly pulsating mass.
The Dunkl-Caldirola-Kanai oscillator, characterized by an exponentially varying mass, effectively models quantum damping effects. Moreover, the harmonic oscillator with a strongly pulsating mass captures periodic energy fluctuations, relevant for systems such as cavity quantum electrodynamics. These results underscore the unique ability of the Dunkl formalism to incorporate reflection symmetries and handle time-dependent parameters, leading to parity-specific propagators and wave functions expressed in terms of modified Bessel functions, instead of the standard trigonometric or exponential functions for the standard case.
Beyond theoretical insights, the practical implications of this work are significant. The Dunkl formalism provides a framework for studying quantum dissipative systems, where exponentially varying mass can model damping effects. Furthermore, it is well-suited for investigating non-Hermitian quantum mechanics, where reflection symmetries could reveal new physical phenomena. The exact solutions derived here could also serve as benchmarks for numerical simulations of time-dependent quantum systems, aiding in the development of more efficient computational techniques.
Future studies could explore the interplay between Dunkl mechanics and non-Hermitian quantum systems, potentially uncovering new dynamical regimes. Additionally, extending the formalism to multi-dimensional systems or PT-symmetric systems would broaden its applicability and relevance.
@@ Acknowledgments
The authors extend their sincere gratitude to the anonymous reviewer for his/her insightful comments and constructive feedback, which have significantly contributed to the improvement of this manuscript. B. C. L. is grateful to Excellence Project PřF UHK 2211/2023-2024 for the financial support.